OpenStudy (mary1999):

consider the translation (x,y)->(x-3,y+6) the preimage of (-7,-4) is

OpenStudy (zarkon):

what have you tried?

OpenStudy (hanamalik):

What is it asking?

OpenStudy (zarkon):

let \[f(x,y)=(x-3,y+6)\] for what values of \(x,y\) does \(f(x,y)=(-7,-4)\)

OpenStudy (mary1999):

I thought it was -7-3 and -4+6

OpenStudy (zarkon):

no...that would be the image of (-7,-4) not the preimage

OpenStudy (mary1999):

so how do i figure the preimage

OpenStudy (zarkon):

solve x-3=-7 and y+6=-4

OpenStudy (mary1999):

x=-4 and y=-10

OpenStudy (zarkon):

put it as an ordered pair (-4,-10)

OpenStudy (mary1999):

so that is the answer

OpenStudy (zarkon):

yes